Jan is a 13-year-old exchange student from Argentina. His father is 44. How many years until he is half his father's age?
Found 2 solutions by macston, Alan3354:
Answer by macston(5194)

.
J=Jan's age=13; F=Father's age=44;
x=years until Jan is half father's age
.
F+x=2(J+x)
44+x=2(13+x)
44+x=26+2x
18=x
ANSWER: In 18 years, Father's age will be twice Jan's age.
.
F+x=2(J+x)
44+18=2(13+18)
62=2(31)
62=62
.

Answer by Alan3354(69443)

Jan is a 13-year-old exchange student from Argentina. His father is 44. How many years until he is half his father's age?
-------------
The difference is 31 years.
--> 31 & 62